OpenStudy (anonymous):

The scores for the first five basketball games were: 88,90,94,98,110 If the score of the sixth game was 98, how is the median changed? increased by 2 decreased by 1 increased by 1 no change weak positive strong negative weak negative no correlation

OpenStudy (anonymous):

increased by 2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

88,90,94,98,110 median=94 88,90,94,98,98,110 median=(94+98)/2=96 96-94=2

OpenStudy (anonymous):

yw:)